OpenStudy (anonymous):

Determine the measure of the acute angle to the nearest degree. tan t°=√3

OpenStudy (anonymous):

tan t = sqrt(3) or tan t = tan 60 hence t=60

OpenStudy (anonymous):

how do you get tan 60?

OpenStudy (anonymous):

or like hjow is t=60?

OpenStudy (anonymous):

take an equilateral triangle now each angle of the triangle is 60

OpenStudy (anonymous):

now from any vertex drop a perpendicular to the opposite side

OpenStudy (anonymous):

let the length of each side be x now the length of the perpendicular so drooped can be found out by using Pythogarass theorem length of perpendicular = sqrt(x^2 -(x/2)^2) = srqt(3) /2 *x now find the tan 60 in that triangle whic equals = (srqt(3) /2 *x )/ x/2 or tan60 =sqrt(3) i hope you got it

OpenStudy (anonymous):

ohhhh i get it now

OpenStudy (anonymous):

wow i feel so dumb

OpenStudy (anonymous):

but thank you!!